Convert the following formula to conjunctive normal form: A ⇒ ( B ∧ C )


A. ( ¬A ∨ B ) ∧ ( ¬A ∨ C )


B. ( A ∧ ¬B ) ∨ ( A ∧ ¬C )


C. ( ¬A ∧ B ) ∨ ( ¬A ∧ C )


D. ( A ∨ ¬B ) ∧ ( A ∨


1
Expert's answer
2022-03-20T14:37:17-0400

"A\,\,is\,\,correct:\\A\rightarrow \left( B\land C \right) =\lnot A\lor \left( B\land C \right) =\\=\left( \lnot A\lor B \right) \land \left( \lnot A\lor C \right)"